# Privacy Technologies for Financial Intelligence

### v1

@@ 4.2 Transaction Flow Tracing and Typology Matching
A key task in financial intelligence supported by transaction-flow tracing is the identification of accounts that match certain money-laundering or terrorism-financing (ML/TF) typologies. The NDIS fraud described in [BILC23] serves as an example. Australia’s National Disability Insurance Scheme (NDIS) provides financial support to eligible participants to improve their quality of life. However, continuous fraudulent activities against the NDIS scheme have resulted in several billions of dollars in losses from its intended use [AUS20]. These fraudulent transactions share common patterns, crossing multiple organisations and eventually being transferred overseas. No single institution in the financial system would have all the data required to identify accounts and entities that exhibit such behaviour. We will show in this section how privacy technologies can help tackle this broad class of problems.
@@
The FinTracer solution proposed in [BILC23] is basically a way to perform an efficient and privacy-preserving breadth-first search across a large distributed graph in homomorphically-encrypted space, allowing Financial Intelligence Units like AUSTRAC to work with law-enforcement agencies and financial institutions to start with some known suspicious entities and find their footprint across the financial system
@@
    without having to disclose the suspicious entities to financial institutions, thus avoiding the risk of accidentally tipping off the entities; and
@@
    without having to violate the privacy of unrelated entities even though their data need to be processed.
The key observation behind this algorithm is that large distributed graphs can be represented using sparse matrices and graph traversal can be expressed as matrix multiplication operations [Bur21]. This basic graph search algorithm can serve as a building block to detect a large range of financial crime typologies (i.e. graph patterns), thus significantly advancing the state-of-the-art in intelligence agency’s ability to detect complex financial crimes in a responsible, privacy-preserving manner at a large scale.
[…]
Figure 7: A simplified transaction network to demonstrate FinTracer on the NDIS example. Accounts with cross organisation transactions are linked by directed edges. There are four financial institutions A, B, C and D, each having a number of accounts. The account a1 in A directly received NDIS payment.
To see how this works, suppose AUSTRAC is interested in identifying the accounts (from multiple financial institutions) that transferred NDIS funds overseas. These accounts either received funds directly from NDIS or indirectly via a limited number of cross-institution transactions. The FinTracer tag $t_{\leq}$ is implemented to keep track of accounts with at most $k$ cross-institution transactions. These tags are homomorphically encrypted to securely follow NDIS funds in the entire transaction network. The key components of the FinTracer process are the ElGamal encryption scheme [ElG85] over the additive Curve25519 for its performance efficiency [Ber06], and the aggregations of the tag values across financial institutions over multiple iterations of the breadth-first search algorithm.
Figure 7 is a simplified example of the NDIS fraud scenario. In this example, there are four financial institutions. The only account that directly received suspicious NDIS fund is a1 in institution A. All institutions are only aware of the source and destination accounts that are involved in the transactions they process. In its simplest form (Algorithms 1 and 2 [BILC23]), the FinTracer protocol does the following steps:
  • Initialisation: Each financial institution sets the initial FinTracer tag of each of its customer account $x$ to either $\text{Enc}(1)$ or $\text{Enc}(0)$, depending on whether $x$ receive NDIS fund directly. In the example above,
@@
    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(2)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(1)$ and $t_{\leq}(d1)=\text{Enc}(0)$.
@@
    • $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(3)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(3)$ and $t_{\leq}(d1)=\text{Enc}(2)$.
@@
An open-source, scalable implementation of FinTracer can be found at http://github.com/AUSTRAC. The implementation includes a C/Cuda library for additive HE with ElGamal over the Ed25519 elliptic curve. A number of algorithms for different graph operations are included in http://github.com/AUSTRAC/ftillite/Documentation/Purgles. The extension of FinTracer, which allows only linear operations because of the use of additive HE, to non-linear set and graph operations to allow the detection of more-or-less arbitrarily complex criminal typologies is described in [BCF23]. Note that differential privacy noise is added in several steps in these algorithms to avoid variations of database-reconstruction attacks [DN03, DMT07].
Finding key players in transaction networks
Once the small number of accounts and entities that matched a certain ML/TF typology is detected using a technology like FinTracer, an FIU like AUSTRAC has the legal power to request much more comprehensive information about those accounts and entities from financial institutions because the request for information or notice to produce information is now targeted and backed by genuine grounds for suspicion. (The remaining small-scale privacy concerns can mostly be addressed using differentially private graph search algorithms like [KRWY16].) With such a dataset, it then makes sense to calculate connected components and risk metrics like PageRank scores and betweenness centrality scores, some of which have been shown to be useful features to enhance the accuracy or confidence of machine learning or rule-based financial intelligence classification models [MCF+17, SvHA+19, vEDvdB+24].
@@ 4.3 Collaborative Model Learning
The use of supervised learning and anomaly detection algorithms to learn financial crime models from data has long been a topic of interest within financial intelligence [SvHA+19, vEDvdB+24], but the practical utility of such approaches has been hampered by a lack of quality labelled data in the case of supervised learning algorithms, and high false positive rates in the case of anomaly detection algorithms.
@@
To enable FL, we would need accounts and transaction data from multiple financial institutions to be matched, often probabilistically, at scale before a machine learning algorithm can be run. In the case of supervised learning, the participating parties would also need to share, as labelled data, entities that are either known or suspected to be engaging in financial crime. This is obviously highly intrusive to entities who are not involved in financial crimes, and also causes the kind of tipping-off issues discussed in Section 4.1 for those (suspected to be) involved in financial crimes. All these imply that we need scalable and secure ways to perform large-scale probabilistic data matching and model learning in a privacy-preserving manner. The publication and reuse of machine learning models by multiple organisations also open the door to attacks like membership inference that need to be managed [MPP+21].
@@
As discussed, the first step in FL is privacy-preserving data matching. Unlike transaction flow tracing, where the source and target of each transaction are uniquely identifiable for exact matching, there can be inconsistencies in how different financial institutions collect and store individual information, such as names, addresses, driver’s licences, business names, and other details. Consequently, data matching and entity resolution in federated learning – both within and across financial institutions – are often performed probabilistically. A comprehensive description of key privacy-preserving data matching/record linkage algorithms, including those that are built on Bloom Filters as described in Section 3.1, can be found in [CRS20].
We now describe a simplified example of federated learning to illustrate its applicability in financial intelligence. The task is to learn a linear regression risk model using probabilistically matched data from multiple participating parties, while preserving privacy in a distributed manner. The target of the regression is the individual risk score, and the features for the regression are derived from all the data held by the REs; the setup is shown in Figure 8 for three parties.
More formally, given a training set $\{({\bf\it x}_{l},y_{l})\}_{l=1,\dots,m}$ that includes $m$ individuals, where the feature set ${\bf\it x}_{l}={\bf\it x}_{l}^{(1)}\cup\cdots\cup{\bf\it x}_{l}^{(p)}$ is the union of the features from all parties, the task is to fit a linear regression model
${\bf\it Y}={\bf\it X}{\bf\it w}+{\bf\it\epsilon}$
that predicts the target variable ${\bf\it Y}$ as accurate as possible.
Figure 8: Federated learning setup for privacy-preserving risk modelling
@@
    The FIU generates a public-private key pair ($pk$, $sk$) and shares the public key $pk$ with the REs.
@@
In the remainder of this subsection, we provide a survey of some innovative use of FL that holds promise for financial intelligence. In the work of [SZK+22], we see a reasonably compelling demonstration of how FL can significantly enhance the detection capabilities of financial crime models. Through a series of case studies and rigorous experimental evaluations, this research highlights the dual achievement of improved detection performance and robust data privacy. The study underscores how FL can elevate the precision of financial crime detection while safeguarding sensitive information.
Building upon this foundation, [DNM+23] introduces a novel approach that advances secure and collaborative model training across multiple financial institutions. By integrating HE with FL, this approach achieves secure model aggregation and parameter updates. The experimental results and evaluations presented reveal the approach’s effectiveness in preserving data privacy and security, thus fostering a collaborative environment for financial data analysis across institutions.
Further extending this narrative, [HHIL+17] presents an end-to-end scheme designed for linear classifiers, employing additive HE to counteract threats from semi-honest adversaries in vertical FL scenarios. This research delves into the impact of entity resolution on learning performance, utilising a Taylor approximation for the loss and gradient functions. The application of Paillier additive HE within this context enables privacy-preserving computations, showcasing an innovative method for handling sensitive data securely.
Additionally, [BP20] proposes a FL scheme tailored for fraud detection. In this scheme, individual financial organisations enhance their model parameters with DP noise before transmitting them to a central server. This strategy effectively diminishes the server’s ability to recover private information, advancing the cause of privacy in fraud detection efforts.
Though these FL techniques have yet to achieve industrial-grade maturity, they collectively offer a glimpse into the forefront of financial intelligence. Their contributions highlight ongoing advancements and set the stage for future innovations in the field.
@@
There are several industrial-grade solutions like [ZCN17, ZNCC18] for conducting probabilistic matching of databases with millions to billions of entities and these can be used in conjunction with techniques like Bloom filters and HE to produce somewhat scalable privacy-preserving probabilistic data-matching algorithms. While federated learning is now a mature technology, progress in using federated learning at the industrial scale for financial intelligence applications has been slow because, in most cases, there is not an abundance of labelled data to support the acquisition of robust and trusted machine-learning models with low false-positive rates. With further advances in FL and a general uplift in the maturity of FIUs in adopting privacy technologies, we expect to see more progress in this area in the future.
5 Discussion and Conclusion

### v2

@@ 4.2 Transaction Flow Tracing and Typology Matching
A key task in financial intelligence supported by transaction-flow tracing is the identification of accounts that match certain money-laundering or terrorism-financing (ML/TF) typologies. The NDIS fraud described in [10] serves as an example. Australia’s National Disability Insurance Scheme (NDIS) provides financial support to eligible participants to improve their quality of life. However, continuous fraudulent activities against the NDIS scheme have resulted in several billions of dollars in losses from its intended use [4]. These fraudulent transactions share common patterns, crossing multiple organisations and eventually being transferred overseas. No single institution in the financial system would have all the data required to identify accounts and entities that exhibit such behaviour. We will show in this section how privacy technologies can help tackle this broad class of problems.
@@
The FinTracer algorithm described in [10] is a computationally efficient way to conduct breadth-first search on large distributed graphs in homomorphically-encrypted space. The algorithm enables Financial Intelligence Units to start with some known suspicious entities and then work with law-enforcement agencies and financial institutions to find the entire footprints of those entities across the financial system
@@
    without having to disclose the actual identities of the suspicious entities to financial institutions, thus avoiding the risk of accidentally tipping off the entities; and
@@
    without having to compromise the privacy of unrelated entities even though their data need to be processed.
The key insight behind the FinTracer algorithm is that large distributed graphs can be represented as sparse matrices and the breadth-first search algorithm can be expressed as a sequence of matrix multiplication operations [12]. The FinTracer algorithm can be used as a foundational routine to detect a wide range of financial crime typologies (i.e. graph patterns). Its adoption can thus meaningfully advance the state-of-the-art in financial intelligence agency’s ability to detect complex financial crimes that span multiple organisations in a privacy-preserving manner at scale.
a1a2a3a4a1b1b2b3b4b5c1c2c3c4c5d1d2d3d4 AA $B$ $C$ $D$
Figure 5: A simplified transaction network to demonstrate FinTracer. Accounts with cross-organisation transactions are linked by directed edges. There are four financial institutions A, B, C and D, each having a number of accounts. The account $a_{1}$ in A directly received NDIS payment.
To see how this works, suppose AUSTRAC is interested in identifying the accounts (from multiple financial institutions) that transferred NDIS funds overseas. These accounts either received funds directly from NDIS or indirectly via a limited number of cross-institution transactions. The FinTracer tag $t_{\leq}$ is implemented to keep track of accounts with at most $k$ cross-institution transactions. These tags are homomorphically encrypted to securely follow NDIS funds in the entire transaction network. The key components of the FinTracer process are the ElGamal encryption scheme [24] over the additive Curve25519 for its performance efficiency [7], and the aggregations of the tag values across financial institutions over multiple iterations of the breadth-first search algorithm.
Figure 5 is a simplified example of the NDIS fraud scenario. In this example, there are four financial institutions. The only account that directly received suspicious NDIS funds is a1 in institution A. All institutions are only aware of the source and destination accounts that are involved in the transactions they process. In its simplest form (Algorithms 1 and 2 [10]), the FinTracer protocol does the following steps:
  • 1.
    Initialisation: Each financial institution sets the initial FinTracer tag of each of its customer account $x$ to either $\text{Enc}(1)$ or $\text{Enc}(0)$, depending on whether $x$ receive NDIS funds directly. In the example above,
    • (a)
@@
    • (a)
      $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(2)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(1)$, $t_{\leq}(d1)=\text{Enc}(0)$.
  • 4.
@@
    • (a)
      $t_{\leq}(a1)=\text{Enc}(1)$, $t_{\leq}(\{b1,b2\})=\text{Enc}(3)$, $t_{\leq}(\{c1,c2\})=\text{Enc}(3)$, $t_{\leq}(d1)=\text{Enc}(2)$.
@@
Finding key players in transaction networks
Once the small number of accounts and entities that matched a certain ML/TF typology is detected using a technology like FinTracer, an FIU like AUSTRAC has the legal power to request much more comprehensive information about those accounts and entities from financial institutions because the request for information or notice to produce information is now targeted and backed by genuine grounds for suspicion. (The remaining small-scale privacy concerns can mostly be addressed using differentially private graph search algorithms like [34].) With such a dataset, it then makes sense to calculate connected components and risk metrics like PageRank scores and betweenness centrality scores, some of which have been shown to be useful features to enhance the accuracy or confidence of machine learning or rule-based financial intelligence classification models [43, 54, 62].
@@
An open-source, scalable implementation of FinTracer can be found at http://github.com/AUSTRAC. The implementation includes a C/Cuda library for additive HE with ElGamal over the Ed25519 elliptic curve. A number of algorithms for different graph operations are included in http://github.com/AUSTRAC/ftillite/Documentation/Purgles. The extension of FinTracer, which allows only linear operations because of the use of additive HE, to non-linear set and graph operations to allow the detection of more-or-less arbitrarily complex criminal typologies is described in [9]. Note that differential privacy noise is added in several steps in these algorithms to avoid variations of database-reconstruction attacks [21, 22].
Techniques described in this section can also be extended to provide more general SQL-like queries on encrypted data, using ideas described in CryptDB [48, 63]. In this area, Enveil and Duality provide industry-grade solutions with their products, Enveil ZeroReveal Search and Duality Query Engine. Such solutions allow organisations to run SQL-like queries on decentralised databases, ensuring data remains within the owner’s environment and supporting compliance with data sovereignty laws. Additionally, these secure systems encrypt both query content and results, keeping them private and visible only to the querier.
@@ 4.3 Collaborative Model Learning
The use of supervised learning and anomaly detection algorithms to learn financial crime models from data has long been a topic of interest within financial intelligence [54, 62], but the practical utility of such approaches has been hampered by a lack of quality labelled data in the case of supervised learning algorithms, and high false positive rates in the case of anomaly detection algorithms.
@@
To enable FL, we would need accounts and transaction data from multiple financial institutions to be matched, often probabilistically, at scale before a machine learning algorithm can be run. In the case of supervised learning, the participating parties would also need to share, as labelled data, entities that are either known or suspected to be engaging in financial crime. This is obviously highly intrusive to entities who are not involved in financial crimes, and also causes the kind of tipping-off issues discussed in Section 4.1 for those (suspected to be) involved in financial crimes. All these imply that we need scalable and secure ways to perform large-scale probabilistic data matching and model learning in a privacy-preserving manner. The publication and reuse of machine learning models by multiple organisations also open the door to attacks like membership inference that need to be managed [44].
@@
As discussed, the first step in FL is privacy-preserving data matching. Unlike transaction flow tracing, where the source and target of each transaction are uniquely identifiable for exact matching, there can be inconsistencies in how different financial institutions collect and store individual information, such as names, addresses, driver’s licences, business names, and other details. Consequently, data matching and entity resolution in federated learning - both within and across financial institutions - are often performed probabilistically. A comprehensive description of key privacy-preserving data matching/record linkage algorithms, including those that are built on Bloom Filters can be found in [18].
We now describe a simplified example of federated learning to illustrate its applicability in financial intelligence. The task is to learn a linear regression risk model using probabilistically matched data from multiple participating parties, while preserving privacy in a distributed manner. The target of the regression is the individual risk score, and the features for the regression are derived from all the data held by the REs. The setup is shown in Figure 6 for three parties.
More formally, given a training set $\{({\bf\it x}_{l},y_{l})\}_{l=1,\dots,m}$ that includes $m$ individuals, where the feature set ${\bf\it x}_{l}={\bf\it x}_{l}^{(1)}\cup\cdots\cup{\bf\it x}_{l}^{(p)}$ is the union of the features from all parties, the task is to fit a linear regression model ${\bf\it Y}={\bf\it X}{\bf\it w}+{\bf\it\epsilon}$ that predicts the target variable $\bf\it Y$} as accurate as possible.
Figure 6: Federated learning setup for privacy-preserving risk modelling using linear regression
@@
    The FIU generates a public-private key pair ($pk$, sksk) and shares the public key $pk$ with the REs.
@@
In the remainder of this subsection, we provide a survey of some innovative use of FL that holds promise for financial intelligence. In the work of [59], we see a reasonably compelling demonstration of how FL can significantly enhance the detection capabilities of financial crime models. Through a series of case studies and rigorous experimental evaluations, this research highlights the dual achievement of improved detection performance and robust data privacy.
Building upon this foundation, [19] introduces a novel approach that advances secure and collaborative model training across multiple financial institutions. By integrating HE with FL, this approach achieves secure model aggregation and parameter updates. The experimental results and evaluations presented reveal the approach’s effectiveness in preserving data privacy and security, thus fostering a collaborative environment for financial data analysis across institutions.
Further extending this narrative, [29] presents an end-to-end scheme designed for linear classifiers, employing additive HE to counter semi-honest adversaries in vertical FL scenarios. This research delves into the impact of entity resolution on learning performance, utilising a Taylor approximation for the loss and gradient functions. The application of Paillier additive HE within this context enables privacy-preserving computations, showcasing an innovative method for handling sensitive data securely.
Additionally, [13] proposes a FL scheme tailored for fraud detection. In this scheme, individual financial organisations enhance their model parameters with DP noise before transmitting them to a central server. This strategy effectively diminishes the server’s ability to recover private information, advancing the cause of privacy in fraud detection efforts.
The above FL techniques collectively offer a glimpse into the forefront of financial intelligence. Their contributions highlight ongoing advancements and set the stage for future innovations in the field.
@@
There are several industrial-grade solutions like [64, 65] for conducting probabilistic matching of databases with millions to billions of entities and these can be used in conjunction with techniques like Bloom filters and HE to produce somewhat scalable privacy-preserving probabilistic data-matching algorithms. Privitar SecureLink [^7] SecureLink website offers a solution for securely joining data from various contributors without revealing sensitive identifiers or directly linking data across organisations.
While federated learning is now a mature technology, progress in using federated learning at the industrial scale for financial intelligence applications is still somewhat hampered by general lack of labelled data to support the acquisition of robust and trusted machine-learning models with low false-positive rates. Having said that, several case studies and investigations have validated and implemented privacy-preserving solutions for specific scenarios. In particular, industrial solutions developed by Enveil, Duality, and Inpher have scaled effectively to address real-world challenges, with some initiatives leading to collaborations with other organisations to create domain specific products. The details are described in Section 5.
5 Real-World Case Studies
In this section, we describe several case studies that have been trialled and, in some cases, deployed in real-world applications. The presentation provides an update on some key projects first listed in [41] and includes more recent additional work; the emphasis is on the business problem, participating financial institutions, and achieved business outcomes (when that is available.) The case studies vary in scope and complexity, with some focusing on a singular use case while others cover all the three major use cases described in Section 4. It is important to note that in the practical context, AML/CTF use cases can be affected by a large number of factors, including computational cost, operational cost, technical complexity, efficiency, data availability, quality and interoperability. These case studies are summarised in Table 1.
Table 1: Overview of several case studies by different jurisdictions.
Project Participants Use case
Enveil ZeroReveal Inter-Bank Secure and Private Data Collaboration [41] An EU-based financial institution and a third-party data provider Transaction flow tracing and typology matching
Duality SecurePlus Query [41] UK and Canadian Banks Transaction flow tracing and typology matching
DANIE consortium [41] UK-based consortium of banks Suspicious persons list
Privitar SecureLink [41] Public and private Institutes in an EU member state Suspicious persons list
BNY Mellon and Inpher XOR [31] BNY Mellon Collaborative learning
UK Tri-bank initiative [42] Three UK banks Transaction flow tracing
AUSTRAC’s Fintel Alliance Alerting Project [14, 58] Led by AUSTRAC Australia and delivered by the public-private partnership Fintel Alliance Transaction flow tracing
AML Bridge Platform [25] Ten banks and several non-banks in Estonia Transaction flow tracing and typology matching
Transaction Monitoring Netherlands (TMNL) [61] Five banks in Netherlands Transaction flow tracing
DeepProtect [32] Five banks in Japan Collaborative learning
Project Aurora [8] Led by Bank for International Settlements and participated by 29 financial institutions in six European countries Collaborative learning
COSMIC [25, 1] Owned and operated by the Monetary Authority of Singapore (MAS). The is a co-development with the Commercial Affairs Department (CAD) of the Singapore Police Force, and six major banks (Citibank, DBS, HSBC, OCBC, Standard Chartered Bank and UOB). Transaction flow tracing and typology matching
Enveil ZeroReveal Inter-Bank Private Collaboration [41]
This is an initiative, started in 2019, by the UK Financial Conduct Authority (FCA) Global AML and Financial Crime TechSprint and Enveil to develop a system that could enable data and knowledge sharing between stakeholders to identify and prevent money flows within and between criminal networks. This initiative was supported by several private UK banks including EY, BAE Systems, Refinitiv, HSBC, Barclays and ING. Their proposed system used homomorphic encryption to provide secure and automated multiparty collaboration which allows banks to query about higher risk customers. Since that FCA TechSprint, the Enveil suite of solutions have developed in maturity and are now used in multiple sensitive domains. [^8] Enveil Wins Army Linchpin Contract to Deliver Secure AI
Duality SecurePlusTM Query on Financial Data [41]
The product Duality SecurePlusTM Query enables financial crime and compliance teams to carry out collaborative investigations across groups, institutions and borders by keeping sensitive data encrypted throughout the process. This product allows cross-border collaboration between financial crime and compliance teams. The product allows participant to query distributed data silos while keeping the sensitive data encrypted throughout the querying process. A case study between several banks in the UK and Canada demonstrated the capability of banks to match customers that includes external data sets, without disclosing their original query to the data owner. This allows different financial intelligence teams to securely query each other’s data. The system also supports to reduce false positives, improve outcomes, increase efficiency, and most importantly effectively expose financial crime networks. [^9] Leverage Secure Data Sharing to Fight Money Laundering
DANIE Consortium [41]
Project DANIE is a collaborative initiative involving multiple financial institutions from EU Member States, aimed at enhancing banking data quality through the utilisation of privacy technologies. The project employs a combination of secure hardware and software components, along with secure multiparty computation and homomorphic encryption. The primary objective of DANIE is to improve data quality by matching datasets across participating organisations to identify discrepancies and outliers, while simultaneously preserving the anonymity and security of the matched data through these technologies. In addition to its application in enhancing data reconciliation and subsequently improving the Know Your Customer (KYC) process, DANIE has the potential to be adapted for addressing the Suspicious Person List (SPL) cases discussed in Section 4.1. This can be achieved by encrypting an SPL and uploading it to the DANIE platform to request confirmation of each individual’s presence within the participating financial institutions.
BNY Mellon and Inpher XOR [31]
The Bank of New York Mellon Corporation uses machine-learning (ML) based fraud detection algorithms to find hidden patterns in money laundering networks. Together with the Inpher’s XOR platform, BNY Mellon built a novel collaborative fraud detection framework leveraging Secure Multiparty Computation. Inpher’s technology enabled BNY Mellon to access sensitive data across borders or jurisdictional silos for training ML algorithms such as linear and logistic regression as well as XGBoost. The solution achieved a high level of precision, computing models identically as if they were computed in clear text. BNY Mellon’s project demonstrated that the number of false positives was reduced by 20% and the number of true positives was increased by 20%. The solution thus reduced investigation workload and increased fraud detection efficiency.
UK Tri-bank initiative [42]
This is a proof-of-concept (POC) system developed by three UK banks to analyse historical data on small and medium sized corporate client datasets. The system used the FutureFlow Cross-Bank Financial Crime Analytics Platform to analyse transaction data between banks. In their POC study, banks analysed approximately 200,000 accounts for money laundering risk. In this system, the transaction data is encrypted to protect sensitive data before using to identify links between the transactions. This allowed the banks to identify potential money laundering networks based on the patterns in the transaction data.
AUSTRAC’s Fintel Alliance Alerting Project [14]
The FINTEL Alliance comprises representatives from both public and private partner organizations that collaborate to address complex or emerging crimes affecting the community, necessitating a joint public-private approach. The aim of this project is to develop different approaches which would allow the identification of complex money transferring networks used by criminal gangs. The project used a federated platform with homomorphic encryption technologies where entities could report or provide access to the agreed dataset via an API. The project was tested in a simulated federated architecture using a sample of real data, covering more than 100 million accounts.
AML Bridge Platform [25]
AML Bridge is a secure platform provided by Salv. The platform uses an end-to end encryption technique which allows financial institutions to securely exchange of pseudonymised transaction data. The platform was first tested as a POC on four of Estonia’s largest banks and later expanded to all Estonian banks. In their POC data owners can also limit what queries it allows from its own data sources while the requesting party is not revealed to the data owner. In their querying protocol, the platform first encrypts the data at the source. Next data is aggregated with other parties’ and stored in three trusted hosts using secure multiparty computation. Finally, within the requesting party’s trusted environment, the platform aggregates the data results without revealing the original data source(s).
Transaction Monitoring Netherlands (TMNL) [61]
This is an initiative consists of five Dutch banks including ABN AMRO, ING, Rabobank, Triodos Bank and De Volksbank. The main aim of this initiative is to combine transaction data of business accounts from these banks to gain a complete view of money flows by criminal networks. The initiative uses a central trusted data store where each bank could upload their data. The data could then aggregate by combining these transactions and the consortium (the group of banks) could monitor and report potentially suspicious patterns to the others. central data store where the consortium combines these transactions, monitors them, and reports potentially suspicious patterns. The central data store encrypts data at rest, which can only be viewed when when suspicions of money laundering or terrorist financing arise Sensitive information is encrypted and viewed only when suspicions of money laundering or terrorist financing arise.
DeepProtect [32]
This is research project by National Institute of Information and Communications Technology (NICT) in Japan. In this project, NICT collaborated with five banks in Japan including Chiba Bank, Ltd., MUFG Bank, Ltd., the Chugoku Bank, Ltd., Sumitomo Mitsui Trust Bank, Ltd., and the Iyo Bank, Ltd. The aim of this project was to build models for detecting criminals’ bank accounts and fraudulent transactions in customers accounts. DeepProtect employs a FL framework which enables each bank to run a local model using its own data while collaboratively train a global model with data from other banks. For privacy preservation, the implementation used a homomorphic encryption scheme in a symmetric key environment.
Project Aurora [8]
This is an initiative led by the Bank for International Settlements (BIS). The aim of this project is to identify money laundering between banks and cross-border financial institutions by analysing the connected payment data. The project uses different privacy and security techniques with ML and network analysis methods for identifying potential suspicious transactions. Phase 1 of the project demonstrated that the combinations of these technologies provides more effective identification of money laundering networks compared to traditional siloed and rules-based approaches. In their experimental setting, consisting 29 financial institutions in six different European countries, the project detected potentially up to 3 times more money laundering involved complex schemes and reduced false positives by up to 80%.
COSMIC [1]
COSMIC is a secure digital platform co-developed by Monetary Authority of Singapore (MAS) and six major commercial banks in Singapore including DBS, OCBC, UOB, Citibank, HSBC and Standard Chartered. Released in 2024, COSMIC allows financial institutions to securely share transaction data and information about suspicious accounts with one another. The initial aim of COSMIC is to provide a easier and efficient data sharing platform for FIs to detect and thereby deter criminal activity. Currently, information sharing in COSMIC is voluntary. The platform currently supports three key financial crime risks, (1) misuse of legal persons, (2) misuse of trade finance for illicit purposes and (3) proliferation financing. COSMIC is owned and operated by the MAS.
The above shortlist of projects is quite compelling and is demonstrative of the increasing global interest in the adoption of maturing privacy technologies by major financial institutions, both public and private, to combat financial crimes.
6 Discussion
Legislative Reform and Privacy Technologies